# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

### 30 Mar 2026 v1

@@ 3.1 Why spectral geometry matters for finite-step Newton–Schulz orthogonalization
Let $\mathbf{G}\in\mathbb{R}^{p\times q}$ have rank $r\geq 1$, with compact SVD $\mathbf{G}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, and suppose $p\leq q$. Fix $\alpha\geq\|\mathbf{G}\|_{2}$, set $\mathbf{X}_{0}=\alpha^{-1}\mathbf{G}$, and define $\mathbf{X}_{k+1}=\left(a\mathbf{I}_{p}+b\mathbf{X}_{k}\mathbf{X}_{k}^{\top}+c(\mathbf{X}_{k}\mathbf{X}_{k}^{\top})^{2}\right)\mathbf{X}_{k}$ for $k\in\mathbb{N}$, where $a=3.4445$, b=4.7750b=-4.7750, and $c=2.0315$ are the coefficients used in Muon’s Newton–Schulz iteration in (Jordan et al., ). Let $\phi(s)=as+bs^{3}+cs^{5}$. Then, for each $k\in\mathbb{N}$, $\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}(s_{1}^{(k)},\ldots,s_{r}^{(k)})\mathbf{V}^{\top},s_{i}^{(0)}=\sigma_{i}/\alpha,s_{i}^{(k+1)}=\phi\bigl(s_{i}^{(k)}\bigr),i\in[r],$ and hence
$\frac{1}{\sqrt{r}}\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k}\sigma_{i}/\alpha\bigr)_{+}^{2}\right)^{1/2}.$
Here $(x)_{+}:=\max\{x,0\}$.
See Appendix A for details.
@@
Under the normalization $\alpha=\|\mathbf{G}\|_{F}$, which is common in practical Newton–Schulz implementations (Jordan et al., ; Liu et al., 2025a), let $p_{i}:=\frac{\sigma_{i}^{2}}{\|\mathbf{G}\|_{F}^{2}},\tau_{i}:=\log_{a}\frac{\|\mathbf{G}\|_{F}}{\sigma_{i}}=\log_{a}\frac{1}{\sqrt{p_{i}}}.$ Then Theorem 3.1 implies
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k-\tau_{i}}\bigr)_{+}^{2}\right)^{1/2}.$
Thus $\tau_{i}$ is precisely the iteration at which the $i$-th singular direction leaves the linear regime. In particular, $\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\tau_{r}=\log_{a}\bigl(\kappa_{2}(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\bigr),\tau_{r}-\tau_{1}=\log_{a}\kappa_{2}(\mathbf{G}).$ Therefore, the stable rank determines when the first singular direction enters the nonlinear regime, whereas the condition number determines the width of the transition window over which the remaining singular directions follow.
Theorem 3.1 shows that broad spectra slow finite-step orthogonalization by delaying contraction along small singular directions. To isolate this effect, Figure 1 uses random Gaussian matrices with controlled shapes and spectral spreads. The top row shows the finite-step error in Theorem 3.1, and the bottom row shows the corresponding condition numbers before and after normalization.
To verify the same mechanism on real training trajectories, Figure 2 uses pre-NS momentum matrices from a 2.6B-token Muon run on LLaMA2-130M Touvron et al. (2023) trained on C4 (Raffel et al., 2020). The top row shows the module-wise median of the error in Theorem 3.1, and the bottom row shows the corresponding mean.
(a) $\times1024$
[…]
(d) $\times4096$
(e) $\times1024$
[…]
(h) $\times4096$
@@
(a) Step 1%
(b) Step 10%
(c) Step 50%
(d) Step 100%
(e) Step 1%
(f) Step 10%
(g) Step 50%
(h) Step 100%
Figure 2: Finite-step orthogonalization error across Newton–Schulz steps at 1%, 10%, 50%, and 100% of training. Top: module-wise median; bottom: module-wise mean; shaded bands denote the 25%–75% range. Two-sided row/column normalization decays fastest.
@@ 3.2 Row/column normalization as a zeroth-order surrogate for whitening
(i) Column/right whitening. Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full column rank, and set $\mathbf{D}_{c}:=\operatorname{diag}(\|\mathbf{M}_{:,1}\|_{2},\ldots,\|\mathbf{M}_{:,q}\|_{2})\succ 0$, $\mathbf{N}_{c}:=\mathbf{M}\mathbf{D}_{c}^{-1}$, and $\mathbf{C}_{c}:=\mathbf{N}_{c}^{\top}\mathbf{N}_{c}-\mathbf{I}_{q}$. If $\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}$ is sufficiently small, then
$\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{M}(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}\bigl(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}\bigr),$
where $\mathbf{L}_{c}$ is the unique solution of $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$ In particular,
@@
(ii) Row/left whitening. Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full row rank, and set $\mathbf{D}_{r}:=\operatorname{diag}(\|\mathbf{M}_{1,:}\|_{2},\ldots,\|\mathbf{M}_{p,:}\|_{2})\succ 0$, $\mathbf{N}_{r}:=\mathbf{D}_{r}^{-1}\mathbf{M}$, and $\mathbf{C}_{r}:=\mathbf{N}_{r}\mathbf{N}_{r}^{\top}-\mathbf{I}_{p}$. If $\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}$ is sufficiently small, then
$\operatorname{Orth}_{\ell}(\mathbf{M})=(\mathbf{M}\mathbf{M}^{\top})^{-1/2}\mathbf{M}=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}\bigr),$
where $\mathbf{L}_{r}$ is the unique solution of $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$ In particular,
@@
Thus row normalization separates marginal rescaling from the genuine whitening step. The column-normalized/right analogue and the corresponding two-sided normalization statements are completely symmetric; see Appendix D.
@@ 3.3 Convergence analysis of the RC and R variants
To analyze the MuonEq family, we focus on the two variants that are most informative theoretically: the two-sided RC map and the default row-normalized R map. RC is the strongest spectral corrector among the three forms but induces a more involved alignment term, whereas R yields the cleanest one-sided geometry and convergence argument. Accordingly, Proposition 3.4 studies the pure RC variant, and Theorem 3.5 provides the main guarantee for the default R variant. The column-normalized C map is retained as the embedding-aligned companion in the method family and empirical ablations.
@@
These assumptions are standard in analyses of adaptive and Muon-type methods (Xie et al., 2024; Chen et al., 2025; Li et al., 2023; Wang et al., 2023; Chang and Yuan, 2025; Huang et al., 2025).
@@
Assume Assumptions 3.1, 3.2, and $3.$ Moreover, assume there exists $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty}$ almost surely. Consider the RC variant (4) with $\lambda=0$ and Nesterov disabled, so that $\hat{\mathbf{M}}_{t}=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t}\mathbf{D}_{c,t}^{-1/2}$ for all $t\in[T]$. Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, and $a:=0.2\sqrt{\max\{m,n\}}$. Suppose further that $\varepsilon\geq\frac{4}{5}G_{\infty}^{2}\max\{m,n\}$, and define $\rho_{r}:=\sqrt{1+\frac{nG_{\infty}^{2}}{\varepsilon}},\rho_{c}:=\sqrt{1+\frac{mG_{\infty}^{2}}{\varepsilon}},\chi_{\varepsilon}:=\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}-\frac{3\rho_{r}\rho_{c}-\rho_{r}-\rho_{c}-1}{4}.$ Then $\rho_{r},\rho_{c}\leq 3/2$ and hence $\chi_{\varepsilon}\geq 1/144$. Consequently,
@@
where $C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$
@@
Assume Assumptions 3.1, 3.2, and $3.$ Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$. Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, $a:=0.2\sqrt{\max\{m,n\}}$, and suppose the trajectory-wise NS5 polar-approximation error $\varepsilon_{\mathrm{ns}}$ defined in Lemma G.2 satisfies $\varepsilon_{\mathrm{ns}}<1$. Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}\bigr)\cdot T^{1/4}},$
where $C_{1}^{\mathrm{ns}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)+aL\Bigl(\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)^{2},C_{2}^{\mathrm{ns}}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}.$
@@
Setting $\varepsilon_{\mathrm{ns}}=0$ recovers the exact-polar theorem for the row-normalized R variant. More generally, compared with the two-sided RC analysis, the descent coefficient is now the fixed positive constant $(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}$, so no auxiliary condition of the form $\chi_{\varepsilon}>0$ is needed.
@@
Equation (5) clarifies the difference between the RC and R variants. The first term is the finite-step orthogonalization error. By Theorem 3.1, it is controlled by the singular spectrum of $S(\mathbf{M})$, so the two-sided RC map can provide a stronger spectral correction. The second term is the bias introduced by preprocessing. Proposition 3.4 shows that the pure RC update preserves descent only under the auxiliary positivity condition $\chi_{\varepsilon}>0$. By contrast, the row-normalized map admits the fixed alignment lower bound
$\left\langle\mathbf{M},\operatorname{Orth}(\mathbf{P}(\mathbf{M})\mathbf{M})\right\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}\|_{F},\qquad\mathbf{P}(\mathbf{M}):=\operatorname{diag}\!\bigl(\operatorname{rowsum}(\mathbf{M}\odot\mathbf{M})\bigr)^{-1/2},$
which is exactly why Theorem 3.5 no longer needs a condition of the form $\chi_{\varepsilon}>0$. Thus RC should be viewed as the stronger two-sided spectral corrector, whereas R is the cleaner one-sided variant and therefore the default choice for the hidden-weight setting considered in this paper. The column-normalized C variant is retained as the embedding-aligned companion and is studied empirically rather than through a separate main theorem.
Remark 3.5.
Theorem 3.5 yields the sublinear stationarity estimate
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}=\mathcal{O}\!\left(\frac{1+\ln T}{T^{1/4}}\right).$
Hence, up to the logarithmic factor and the explicit problem-dependent constants, the default R variant has the same $T^{-1/4}$-type baseline dependence as the currently known convergence theory for standard stochastic nonconvex Muon-type methods (Chen et al., 2025; Chang et al., 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025). Improvements beyond this baseline typically require additional mechanisms such as variance reduction (Liu et al., 2025b; Chang et al., 2025; Huang et al., 2025; Qian et al., 2025; Fang et al., 2018; Zhou et al., 2020; Cutkosky and Orabona, 2019; Huang et al., 2021).

### 10 May 2026 v2

@@ 3.1 Why spectral geometry matters for finite-step Newton–Schulz orthogonalization
Let $\mathbf{G}\in\mathbb{R}^{p\times q}$ have rank $r\geq 1$ and compact SVD $\mathbf{G}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ with $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$ and $p\leq q$. Consider the Newton–Schulz iteration
$\mathbf{X}_{0}=\frac{\mathbf{G}}{\|\mathbf{G}\|_{F}},\quad\mathbf{X}_{k+1}=\left(a\mathbf{I}_{p}+b\mathbf{X}_{k}\mathbf{X}_{k}^{\top}+c(\mathbf{X}_{k}\mathbf{X}_{k}^{\top})^{2}\right)\mathbf{X}_{k},$
where $\phi(s):=as+bs^{3}+cs^{5}$ and $q(t):=a+bt+ct^{2}$ satisfy $a>1,0<q(t)\leq a,t\in[0,1].$ Such coefficient choices include those used in (Jordan et al., ; Gyu Yeol Kim, 2026). Then, for each $k\geq 0$,
$\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}\bigl(s_{1}^{(k)},\ldots,s_{r}^{(k)}\bigr)\mathbf{V}^{\top},\qquad s_{i}^{(0)}=\frac{\sigma_{i}}{\|\mathbf{G}\|_{F}},\qquad s_{i}^{(k+1)}=\phi\bigl(s_{i}^{(k)}\bigr).$
Moreover,
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\left(1-\frac{a^{k}}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}}\right)_{+}^{2}\right)^{1/2}.$
Equivalently, the $i$-th singular direction leaves the linear regime at $\tau_{i}=\log_{a}\left(\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}\right)$ . In particular, $\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\tau_{r}=\log_{a}\left(\kappa(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\right),\tau_{r}-\tau_{1}=\log_{a}\kappa(\mathbf{G})$ . See Appendix B for details.
@@
Theorem 3.1 shows that the stable rank controls the onset of the linear to nonlinear transition, while the condition number controls its width across singular directions.
Refer to captionRefer to caption
((a))
Refer to captionRefer to caption
((b))
Refer to captionRefer to caption
((c))
Refer to captionRefer to caption
((d))
@@
To isolate this effect, Figure 1 uses random Gaussian matrices with controlled shapes and spectral spreads. The top row shows the finite-step error in Theorem 3.1, and the bottom row shows the corresponding condition numbers before and after normalization.
To verify the same mechanism on real training trajectories, Figure 2 uses pre-NS momentum matrices from a 2.6B-token Muon run on LLaMA2-130M Touvron et al. (2023) trained on C4 (Raffel et al., 2020). The top row shows the module-wise median of the error in Theorem 3.1, and the bottom row shows the corresponding mean. Appendix J provides a more detailed module-wise view of the same phenomenon: Figure 6 reports singular-value entropy and stable-rank diagnostics for the pre-NS momentum matrices, showing that equilibration reshapes the spectrum seen by finite-step Newton–Schulz.
Refer to captionRefer to caption
((a))
Refer to captionRefer to caption
((b))
Refer to captionRefer to caption
((c))
Refer to captionRefer to caption
((d))
Figure 2: Finite-step orthogonalization error across Newton–Schulz steps at 1%, 10%, 50%, and 100% of training. In each column, the top panel shows the module-wise median and the bottom panel shows the module-wise mean; shaded bands denote the 25%–75% range. Two-sided row/column normalization decays fastest.
@@ 3.2 Row/column normalization as a zeroth-order surrogate for whitening
(i) Column/right whitening.
Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full column rank, and set $\mathbf{D}_{c}:=\operatorname{diag}(\|\mathbf{M}_{:,1}\|_{2},\ldots,\|\mathbf{M}_{:,q}\|_{2})\succ 0$, $\mathbf{N}_{c}:=\mathbf{M}\mathbf{D}_{c}^{-1}$, and $\mathbf{C}_{c}:=\mathbf{N}_{c}^{\top}\mathbf{N}_{c}-\mathbf{I}_{q}$. If $\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{r}(\mathbf{M}):=\mathbf{M}(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}\bigl(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}\bigr),$ where $\mathbf{L}_{c}$ is the unique solution of $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$ In particular,
@@
(ii) Row/left whitening.
Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full row rank, and set $\mathbf{D}_{r}:=\operatorname{diag}(\|\mathbf{M}_{1,:}\|_{2},\ldots,\|\mathbf{M}_{p,:}\|_{2})\succ 0$, $\mathbf{N}_{r}:=\mathbf{D}_{r}^{-1}\mathbf{M}$, and $\mathbf{C}_{r}:=\mathbf{N}_{r}\mathbf{N}_{r}^{\top}-\mathbf{I}_{p}$. If $\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{\ell}(\mathbf{M}):=(\mathbf{M}\mathbf{M}^{\top})^{-1/2}\mathbf{M}=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}\bigr),$ where $\mathbf{L}_{r}$ is the unique solution of $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$ In particular,
@@
Thus row normalization separates marginal rescaling from the genuine whitening step. The column-normalized/right analogue and the corresponding two-sided normalization statements are completely symmetric; see Appendix E.
@@ 3.3 Convergence analysis of the RC and R variants
To analyze the MuonEq family, we focus on the two variants that are most informative theoretically: RC and the default row-normalized R. RC gives stronger two-sided spectral correction but leads to a more involved alignment term, while R has a cleaner one-sided geometry. Accordingly, Proposition 3.4 studies RC. For R, Theorem 3.5 gives the main guarantee for the exact-polar update, and Corollary 3.6 extends it to the finite-step NS5 implementation. Compared with recent Muon analyses Chang et al. (2025); Sato et al. (2025); Shen et al. (2025); Gyu Yeol Kim (2026); Shulgin et al. (2025); Nagashima and Iiduka (2026), our contribution is not to improve the asymptotic rate, but to establish a guarantee closer to the optimizer setting used in practice: the same $\widetilde{\mathcal{O}}(T^{-1/4})$ stationarity rate is proved while explicitly covering decoupled weight decay, a learning-rate schedule independent of the training horizon, and finite-step NS5 through the inexactness constant $\varepsilon_{\mathrm{ns}}$.
@@
These assumptions are standard in analyses of adaptive and Muon-type methods (Xie et al., 2024; Chen et al., 2025; Li et al., 2023; Wang et al., 2023; Chang and Yuan, 2026; Li et al., 2026; Huang et al., 2025).
@@
Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, and $a:=0.2\sqrt{\max\{m,n\}}$. Assume Assumptions 3.1, 3.2, and 3.3. Moreover, assume there exists $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty}$ almost surely. For Algorithm 1, consider the exact-polar RC variant in (4), with Nesterov momentum disabled and $\lambda_{t}=0$. Suppose further that $\varepsilon\geq\frac{4}{5}G_{\infty}^{2}\max\{m,n\}$, and define $\rho_{r}:=\sqrt{1+\frac{nG_{\infty}^{2}}{\varepsilon}},\rho_{c}:=\sqrt{1+\frac{mG_{\infty}^{2}}{\varepsilon}},\chi_{\varepsilon}:=\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}-\frac{3\rho_{r}\rho_{c}-\rho_{r}-\rho_{c}-1}{4}.$ Then $\rho_{r},\rho_{c}\leq 3/2$ and hence $\chi_{\varepsilon}\geq 1/144$. Consequently,
@@
where $C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$ See Appendix G for details.
@@
Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, $\varepsilon=0$ and $a:=0.2\sqrt{\max\{m,n\}}$. Assume Assumptions 3.1, 3.2, and 3.3. For Algorithm 1, consider the exact-polar R variant in (4), with Nesterov momentum disabled and $0\leq\rho<\frac{a}{\sqrt{m}},0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a\sqrt{n}}\,t^{-1/4}$. Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ep}}(1+\ln T)+C_{2,\rho}^{\mathrm{ep}}}{\bigl(a/\sqrt{m}-\rho\bigr)\cdot T^{1/4}},$
where $C_{1,\rho}^{\mathrm{ep}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}(a\sqrt{n}+\rho)^{2}+\sigma^{2}\Bigr)+aL\Bigl(\frac{1}{\sqrt{m}}+\sqrt{n}\Bigr)^{2},C_{2,\rho}^{\mathrm{ep}}=\frac{a\sigma^{2}}{L}+\frac{3L}{2}(a\sqrt{n}+\rho)^{2}.$
Corollary 3.6.
Under the assumptions of Theorem 3.5, replace the exact-polar step by the finite-step NS5 step $\mathbf{O}_{t}:=\mathrm{NS5}(\hat{\mathbf{M}}_{t}),$ where $\mathrm{NS5}$ is the five-step pre-scaled Newton–Schulz map in Lemma H.2. For the fixed horizon $T$, we define
$\varepsilon_{\mathrm{ns}}:=\varepsilon_{\mathrm{ns},T}=\max_{1\leq t\leq T}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}.$
By Lemma H.2, $\varepsilon_{\mathrm{ns}}<1$. Replace the decoupled weight decay condition by $0\leq\rho<\frac{a(1-\varepsilon_{\mathrm{ns}})}{\sqrt{m}},0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}\,t^{-1/4}.$ Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}}{\bigl(a(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}-\rho\bigr)\cdot T^{1/4}},$
where $C_{1,\rho}^{\mathrm{ns}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}+\sigma^{2}\Bigr)+aL\Bigl(\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)^{2}$ , $C_{2,\rho}^{\mathrm{ns}}=\frac{a\sigma^{2}}{L}+\frac{3L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}$. See Appendix I for details.
@@
When $\varepsilon=0$, $\mathbf{D}_{r,t}^{-1/2}$ is interpreted as $(\mathbf{D}_{r,t}^{1/2})^{\dagger}$, leaving zero rows unchanged. Theorem 3.5 is the main exact-polar guarantee for the default R variant. Corollary 3.6 transfers the same $T^{-1/4}$ dependence to the NS5 implementation, up to the explicit $(1\pm\varepsilon_{\mathrm{ns}})$ constants. In Algorithm 1, enabling Nesterov momentum simply amounts to replacing $\mathbf{M}_{t}$ with $\widetilde{\mathbf{M}}t$ in both $\mathbf{D}{r,t}$ and $\hat{\mathbf{M}}_{t}$; this should not change the stated $T^{-1/4}$ stationarity result, consistent with Chen et al. (2025); Chang et al. (2025).
@@
Equation (5) separates finite-step orthogonalization error from preprocessing bias. RC is the stronger two-sided spectral corrector, but Proposition 3.4 requires the auxiliary condition $\chi_{\varepsilon}>0$. By contrast, the row-normalized map satisfies
$\left\langle\mathbf{M},\operatorname{Orth}(\mathbf{P}(\mathbf{M})\mathbf{M})\right\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}\|_{F},\qquad\mathbf{P}(\mathbf{M}):=\bigl(\operatorname{diag}(\operatorname{rowsum}(\mathbf{M}\odot\mathbf{M}))^{1/2}\bigr)^{\dagger},$
so Theorem 3.5 needs no $\chi_{\varepsilon}$-type condition, and Corollary 3.6 only degrades this coefficient by the explicit factor $1-\varepsilon_{\mathrm{ns}}$. Hence the default R variant obeys
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}=\mathcal{O}\left(\frac{1+\ln T}{T^{1/4}}\right),$
matching the standard Muon-type baseline up to logarithmic (Chen et al., 2025; Chang et al., 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025) and constant factors, while C is retained as the embedding-aligned companion and studied empirically (Xu et al., 2026b). Improvements beyond this baseline typically require additional mechanisms such as variance reduction (Yuan et al., 2024; Chang et al., 2025; Liu et al., 2025b; Qian et al., 2025; Fang et al., 2018; Zhou et al., 2020; Cutkosky and Orabona, 2019; Huang et al., 2021).